Solve for k in k-3/2=k+1/5
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

k-3/2=k+1/5 No solution exists

Answer by MathTherapy(10552)


Solve for k in k-3/2=k+1/5
TOO AMBIGUOUS!!
They say "parentheses are free." Why don't you use some?